How do I figure out this problem?
x-5y=2
2x-9y=5

Answer by rfer(16322)

2(5y+2)-9y=5
10y+4-9y=5
y=5-4
y=1
---------------
2x-9*1=5
2x-9=5
2x=14
x=14/2
x=7
--------------
7-5*1=2